# An Imitative Reinforcement Learning Framework for Autonomous Dogfight

### v1

@@ Abstract.
Unmanned Combat Aerial Vehicle (UCAV) dogfight, which refers to a fight between two or more UCAVs usually at close quarters, plays a decisive role on the aerial battlefields. With the evolution of artificial intelligence, dogfight progressively transits towards intelligent and autonomous modes. However, the development of autonomous dogfight policy learning is hindered by challenges such as weak exploration capabilities, low learning efficiency, and unrealistic simulated environments. To overcome these challenges, this paper proposes a novel imitative reinforcement learning framework, which efficiently leverages expert data while enabling autonomous exploration. The proposed framework not only enhances learning efficiency through expert imitation, but also ensures adaptability to dynamic environments via autonomous exploration with reinforcement learning. Therefore, the proposed framework can learn a successful dogfight policy of ‘pursuit-lock-launch’ for UCAVs. To support data-driven learning, we establish a dogfight environment based on the Harfang3D sandbox, where we conduct extensive experiments. The results indicate that the proposed framework excels in multistage dogfight, significantly outperforms state-of-the-art reinforcement learning and imitation learning methods. Thanks to the ability of imitating experts and autonomous exploration, our framework can quickly learn the critical knowledge in complex aerial combat tasks, achieving up to a $100\%$ success rate and demonstrating excellent robustness.
@@ 1. Introduction
In recent years, emerging artificial intelligence technologies such as deep learning have made remarkable advancements in the military field. The increasing significance of artificial intelligence applications in military operations has propelled modern warfare toward intelligent development (Van’t Wout et al., 2019; Benaskeur et al., 2010). Unmanned combat aerial vehicles (UCAVs), as key elements in aerial combat, have become essential in modern military and defense, demonstrating outstanding combat performance in tasks such as aerial strikes, reconnaissance, electronic warfare, and target tracking. The autonomous control of UCAVs has gradually become a primary research focus (Hu et al., 2022; Chamola et al., 2021; Bell and Waag, 2017), and dogfight is one of the most important tasks for UCAVs, which refers to a fight between two or more UCAVs usually at close quarters. This paper mainly focuses on autonomous control policy learning for dogfight tasks between two UCAVs.
@@
There has been substantial research on autonomous control policy learning for UCAV dogfight engagements (Ruan et al., 2022; Ramírez López and Żbikowski, 2018; Trsek, 2007), and most existing works rely on supervised learning and offline imitation learning techniques (Zhang and Huang, 2020; Li et al., 2022; Sandström et al., 2022; Kong et al., 2020). With a delicately designed dataset, these methods can learn a well-performed control policy for the UCAVs. However, the absence of online interactions can lead to the overfitting to these static datasets. In realistic combat environments, opponent policies are everchanging, and unforeseen obstacles may appear suddenly. The UCAV control policies learned with a fixed dataset can hardly handle such dynamic environments. Furthermore, the cost of constructing such a training dataset is expensive as well, and the performance of the learned policy heavily depends on the quality of the dataset. Consequently, it is of great importance to develop more adaptive learning frameworks for UCAV control policy.
@@
Recently, more researchers have investigated reinforcement learning (RL) (Sutton and Barto, 2018) techniques to achieve autonomous policy learning for UCAVs (Kim et al., 2003; Wang and Wei, 2022; Fu et al., 2022; Li et al., 2023; Fan et al., 2022; Cao et al., 2023). Different from supervised learning methods, RL enables the control agent for the UCAV to gradually learn policies through online interactions with the environment. Considering the dynamic property of the dogfight environment, RL is suitable for such a learning problem (Piao et al., 2024), as it supports online exploration and therefore can better adapt to dynamic changes in the environments. However, dogfight policy learning is quite a complex problem, which involves multiple control dimensions, e.g., the direction of flight, target lock, and missile launch. Due to the trial-and-error nature of RL agents, the sample efficiency of RL algorithms in such complex learning scenarios is low, leading to poor success rates of dogfight engagements.
@@
Besides the learning frameworks, the simulated learning environment is substantially important as well. Since directly conducting dogfight policy learning in real environments is impossible, an effective and high-fidelity simulator is urgently required. However, the simulated environments in previous works suffer from the oversimplified action spaces and task setting. For example, the action space in (Fan et al., 2022) is simplified into a discrete predefined maneuver library; (Fu et al., 2022) confines the action space to a two-dimensional simulation environment; (Wang and Wei, 2022; Li et al., 2023; Cao et al., 2023) merely use part of the parameters of the UCAV flight dynamics equations as the action space. As for the task setting, (Kim et al., 2003) solely focuses on the flight control of UCAV without launching missiles; (Wu et al., 2023) tackles missile evasion of UCAV without flight control; (Hsu et al., 2018; Choi et al., 2019; Zhu et al., 2017; Kahn et al., 2018) aim to achieve the objectives of route planning; (Haarnoja et al., 2018; Fu et al., 2022) aim to accomplish the pursuit subtask; (Li et al., 2023) completes the locking of the opponent based on rule-based pursuing. Nevertheless, to successfully shoot down the opponent in a realistic dogfight task, UCAV needs to accomplish three subtasks at least. These include swiftly pursuing the opponent through the optimal path, locking onto the target within a proper range, and launching missiles once the opponent is locked.
@@
To address the aforementioned challenges in dogfight policy learning, we introduce a novel imitative reinforcement learning framework that integrates imitation learning with autonomous exploration by the RL agent. Imitating expert data improves learning efficiency, and autonomous exploration with the RL agent helps adapt to dynamic environments. In this work, the UCAV dogfight problem is formulated as a Markov Decision Process (MDP). To alleviate the limitations of previous dogfight simulators, we construct a highly realistic UCAV simulation environment based on the Harfang3D sandbox (Özbek et al., 2022). The constructed learning environment is endowed with an action space design with high fidelity, controlling UCAV directional rudders, elevators, and ailerons, and thus more realistic than simplified dynamics equation parameter control in previous works. As a replication of the real dogfight process, the simulated dogfight task consists of three stages, i.e., pursuit, lock, and launch, providing a more complex replication of real dogfight processes than previous single-stage works. As an evaluation, we conduct experiments in the constructed dogfight environment. The experiment results demonstrate that the training environment constructed in this paper has high fidelity, effectively simulating the UCAV dogfight process, with a physical model that matches the real-world scenarios better. The proposed learning framework performs excellently in the UCAV multi-dimensional decision-making task, significantly outperforming state-of-the-art deep RL and imitation learning algorithms, achieving higher learning efficiency, more effective air combat policies, and higher success rates. The contributions of this work can be summarized as follows.
@@ 2. Preliminaries
In this section, we review the preliminary knowledge about Harfang3D Sandbox simulator, Markov decision process, and the actor-critic framework.
@@ 2.1. Harfang3D Sandbox
We establish a simulation platform based on the Harfang3D sandbox (Özbek et al., 2022), which is a customizable and flexible tool for combat aircraft dogfight. This platform facilitates the manipulation and access to flight dynamics models and environment states. The sandbox provides simulated aircrafts of five different models, each equipped with unique missile configurations, and a limited missile supply. Therefore, the aircraft must lock onto a target before launching a missile for avoiding missing the target and consequently losing a missile, which increases the task difficulty. Within this simulation environment, information about environment states can be obtained, including aircraft position, attitude, heading angle, and health as well as target lock status, thrust, lock angle, and missile loading. Specially, the sandbox also offers various unmanned aircraft control methods for components including rudders, elevators, ailerons, missile launch, engine thrust, landing gear, and flaps. As illustrated in Figure 1, rudders control the aircraft heading, elevators control the aircraft pitch, and ailerons control the aircraft roll. These three controls enable basic transformations of the aircraft’s attitude.
@@ 2.2. Markov Decision Process
The Markov Decision Process (MDP) (Garcia and Rachelson, 2013) is the theoretical foundation for RL, which provides a formal description of the learning environment of the agent. In this paper, we model the UCAV dogfight environment as an MDP. The MDP is represented as a tuple $(S,A,P,R,\gamma)$, where $S$ is the state space of the environment, $A$ is the action space of the agent, $P$ is the state transition probability, with $P(s_{t+1}|s_{t},a_{t})$ indicating the probability of transitioning from state $s_{t}$ to state $s_{t+1}$ when taking action $a_{t}$ at time $t$, $R$ is the reward function, representing the immediate reward obtained by the intelligent agent after taking action $a$, and $\gamma$ is the reward discount factor, with $\gamma\in[0,1)$ and a $\gamma$ value closer to 1 indicating that the intelligent agent focuses more on long-term rewards. Reinforcement learning aims to learn policy function $\pi:S\rightarrow A$ that maximizes discounted reward $\mathbb{E}_{P,\pi}[\sum_{t=1}^{T}\gamma^{t}R(s_{t},a_{t})]$, where $T$ is the total timestep in an episode.
@@ 2.3. Actor-Critic Framework
The actor–critic framework (Konda and Tsitsiklis, 1999) is a policy optimization paradigm in RL that simultaneously optimizes both the value and policy functions, exhibiting good learning performance in sequential decision problems with continuous action spaces. The actor represents the policy $\pi$ of the agent, which specifies the actions to take at each state. It aims to learn an optimal policy that maximizes the expected cumulative reward. During training, the actor adjusts its policy based on feedback from the critic, gradually improving the policy performance. The critic estimates the value of each action to assess the quality of the policy and assists the actor in refining its policy. The actor–critic framework is based on policy gradient methods, combining back-propagation with value function methods to enhance the policy performance. Based on the actor-critic framework, the agent can learn a policy and corresponding value function during interaction with the environment, enabling high-level decision-making and intelligent behavior.
@@ 3. Related Work
The previous works deal with the UCAV dogfight problem either with the conventional methods or using the advanced approaches with deep neural networks.
3.1. Conventional Methods
Before deep learning methods are employed in the dogfight policy learning domain, conventional methods are adopted to address tactical decision-making problems in UCAV operation, including expert systems, game theory, and influence diagrams.
Expert knowledge-based methods (Wang et al., 2019; Burgin and Eggleston, 1976; Burgin et al., 1975) involve creating an expert system that infers decisions by directly applying predefined rules within the system. As the rules in the expert system are fixed, the system cannot provide effective control policies when facing situations not covered by the predefined rules. Fu et al. (Fu et al., 2022) proposed an expert system control method based on the rolling time domain to improve the system response. In cases where traditional expert systems fail, their method allows for rapid autonomous decision-making to achieve attack conditions.
Game-theory-based methods (Park et al., 2016) allow modeling aerial combat scenarios as game problems by treating UCAV as players in the game. The actions of UCAV constitute the policy space, and reward functions are defined based on parameters such as the positions of UCAVs from the two sides. Park et al. (Park et al., 2016) applied differential game theory to generate tactical action combinations for UCAVs. Through a layered decision structure and calculation of a scoring matrix based on differential game theory, optimal actions to deal with dynamic combat environments are determined.
Influence diagram-based methods (Lin et al., 2007) allow the abstraction of the decision network in UCAV operation. These methods involve scenario analysis based on the constructed influence diagram to select the optimal course of action. Zhong et al. (Lin et al., 2007) used multistage influence diagrams to model one-on-one aerial combat scenarios. They transformed the multistage influence diagram into a two-layer optimization problem as a solution to obtain the optimal attack trajectory.
Conventional methods typically rely on predefined rules or models. Because these rules and models are often manually designed, UCAVs lack mechanisms for autonomous learning and adaptation. Consequently, these methods are generally inadequate for dealing with uncertainties on the battlefield.
3.2. Deep Learning Methods
@@
With the rapid development of deep learning, the control for UCAVs has become moreautonomous. Deep learning enables UCAVs to both execute fixed programs and exhibit learning capabilities. In recent years, deep learning techniques such as supervised learning, imitation learning, and reinforcement learning have been widely used to address tactical decision-making problems in UCAV dogfights.
@@
Supervised learning is predominantly used in UCAV dogfight to predict the states after executing actions, providing a basis for action selection. Zhang and Huang (Zhang and Huang, 2020) proposed a deep learning method integrating state evaluation and action selection for making flight action decisions. The state evaluation function scores the next state after executing an action and selects the action that generates the highest score. This approach enables autonomous flight control of UCAVs in one-on-one aerial combat. Li et al. (Li et al., 2022) trained a long short-term memory network with time-series data, mapping historical state sequences onto flight control parameters. They used a stacked sparse autoencoder to reduce data redundancy and speed up decision-making.
@@
Imitation learning (Sandström et al., 2022) is a form of supervised learning method where the model at-tempts to learn and replicate behaviors from examples or expert demonstrations. Imitation learning methods include 1) behavior cloning (BC), 2) inverse reinforcement learning, and 3) generative adversarial imitation learning.
@@
1. BC (Sandström et al., 2022) involves training a model on expert demonstration data to resemble the expert behaviors. Sandström et al. (Sandström et al., 2022) adopted BC using control signals from human pilots to train a feedforward neural network, enabling the model to mimic the behaviors of human pilots for aircraft flight.
@@
2. Inverse reinforcement learning (Kong et al., 2020) estimates the reward function from expert demonstrations and uses it to guide the training of intelligent agents in reinforcement learning. By extracting the hidden reward function from expert data, the system can understand task objectives from expert demonstrations. Kong et al. (Kong et al., 2020) employ a reward shaping method to address the problem of sparse rewards in reinforcement learning and use maximum entropy inverse reinforcement learning to obtain a shaped reward function, thereby accelerating convergence in training.
@@
3. Generative adversarial imitation learning (Wang and Wei, 2022) uses adversarial generative networks to learn hidden policies from expert demonstrations, imitating expert actions. Wang and Wei (Wang and Wei, 2022) applied this technique to address the problem of sparse rewards in reinforcement learning. They designed internal rewards to stimulate intelligent agents to explore autonomously, thus expediting training.

### v2

@@ Abstract.
Unmanned Combat Aerial Vehicle (UCAV) dogfight, which refers to a fight between two or more UCAVs usually at close quarters, plays a decisive role on the aerial battlefields. With the evolution of artificial intelligence, dogfighting progressively transits towards intelligent and autonomous modes. However, the development of autonomous dogfight policy learning is hindered by challenges such as weak exploration capabilities, low learning efficiency, and unrealistic simulated environments. To overcome these challenges, this paper proposes a novel imitative reinforcement learning framework, which efficiently leverages expert data while enabling autonomous exploration. The proposed framework not only enhances learning efficiency through expert imitation, but also ensures adaptability to dynamic environments via autonomous exploration with reinforcement learning. Therefore, the proposed framework can learn a successful dogfight policy of ‘pursuit-lock-launch’ for UCAVs. To support data-driven learning, we establish a dogfight environment based on the Harfang3D sandbox, where we conduct extensive experiments. The results indicate that the proposed framework excels in multistage dogfighting, and significantly outperforms state-of-the-art reinforcement learning and imitation learning methods. Thanks to the ability of imitating experts and autonomous exploration, our framework can quickly learn the critical knowledge in complex aerial combat tasks, achieving up to a $100\%$ success rate and demonstrating excellent robustness.
@@ 1. Introduction
In recent years, emerging artificial intelligence technologies such as deep learning have made remarkable advancements in the military field. The increasing significance of artificial intelligence applications in military operations has propelled modern warfare toward intelligent development (van2019framework, ; benaskeur2010corals, ). Unmanned combat aerial vehicles (UCAVs), as key elements in aerial combat, have become essential in modern military and defense, demonstrating outstanding combat performance in tasks such as aerial strikes, reconnaissance, electronic warfare, and target tracking. Autonomous control of UCAVs has gradually become a primary research focus (hu2022autonomous, ; chamola2021comprehensive, ; bell2017evaluating, ), with dogfight is one of the most important tasks for UCAVs, which refers to a fight between two or more UCAVs usually at close quarters. This paper mainly focuses on autonomous control policy learning for dogfight tasks between two UCAVs.
@@
There has been substantial research on autonomous control policy learning for UCAV dogfight engagements (ruan2022autonomous, ; ramirez2018effectiveness, ; trsek2007last, ), and most existing works rely on supervised learning and offline imitation learning techniques (zhang2020maneuver, ; li2022decision, ; sandstrom2022fighter, ; kong2020uav, ). With a delicately designed dataset, these methods can learn a well-performed control policy for the UCAVs. However, the absence of online interactions can lead to the overfitting of these static datasets. In realistic combat environments, opponent policies are everchanging, and unforeseen obstacles may appear suddenly. The UCAV control policies learned with a fixed dataset can hardly handle such dynamic environments. Furthermore, the cost of constructing such a training dataset is expensive as well, and the performance of the learned policy heavily depends on the quality of the dataset. Consequently, it is of great importance to develop more adaptive learning frameworks for UCAV control policy.
@@
Recently, more researchers have investigated reinforcement learning (RL) (sutton2018reinforcement, ) techniques to achieve autonomous policy learning for UCAVs (kim2003autonomous, ; wang2022research, ; fu2022uav, ; li2023manoeuvre, ; fan2022air, ; cao2023autonomous, ). Different from supervised learning methods, RL enables the control agent for the UCAV to gradually learn policies through online interactions with the environment. Considering the dynamic property of the dogfight environment, RL is suitable for such a learning problem (piao2024discovering, ), as it supports online exploration and therefore can better adapt to dynamic changes in the environments. However, dogfight policy learning is quite a complex problem, which involves multiple control dimensions, e.g., the direction of flight, target lock, and missile launch. Due to the trial-and-error nature of RL agents, the sample efficiency of RL algorithms in such complex learning scenarios is low, leading to poor success rates of dogfight engagements.
@@
Besides the learning frameworks, the simulated learning environment is substantially important as well. Since directly conducting dogfight policy learning in real environments is impossible, an effective and high-fidelity simulator is urgently required. However, the simulated environments in previous works suffer from the oversimplified action spaces and task setting. For example, the action space in (fan2022air, ) is simplified into a discrete predefined maneuver library; (fu2022uav, ) confines the action space to a two-dimensional simulation environment; (wang2022research, ; li2023manoeuvre, ; cao2023autonomous, ) merely use part of the parameters of the UCAV flight dynamics equations as the action space. As for the task setting, (kim2003autonomous, ) solely focuses on the flight control of UCAV without launching missiles; (wu2023intelligent, ) tackles missile evasion of UCAV without flight control; (hsu2018distributed, ; choi2019deep, ; zhu2017target, ; kahn2018self, ) aim to achieve the objectives of route planning; (haarnoja2018soft, ; fu2022uav, ) aim to accomplish the pursuit subtask; (li2023manoeuvre, ) completes the locking of the opponent based on rule-based pursuing. The flight dynamics model and task setting in (berndt2004jsbsim, ) is realistic, but this simulator is implemented with C++, which is hard to adapt to the deep learning algorithms implemented with Python. To successfully shoot down the opponent in a realistic dogfight task, UCAV needs to accomplish three subtasks at least. These include swiftly pursuing the opponent through the optimal path, locking onto the target within a proper range, and launching missiles once the opponent is locked.
@@
To address the aforementioned challenges in dogfight policy learning, we introduce a novel imitative reinforcement learning framework that integrates imitation learning with autonomous exploration by the RL agent. By imitating expert data, learning efficiency is improved, while autonomous exploration by the RL agent facilitates adaptation to dynamic environments. In this work, the UCAV dogfight problem is formulated as a Markov Decision Process (MDP). To alleviate the limitations of previous dogfight simulators, we construct a realistic UCAV simulation environment based on the Harfang3D sandbox (ozbek2022harfang3d, ). The constructed learning environment features a high-fidelity action space design that enables control over UCAV directional rudders, elevators, and ailerons, making it more realistic than prior works that employed simplified dynamics equation parameters for control. Beyond the action space, the dynamic model in the constructed simulator is also of high fidelity, which is built according to the real flight dynamics models of various aircraft. As a replication of the real dogfight process, the simulated dogfight task consists of three stages, i.e., pursuit, lock, and launch, providing a more complex replication of real dogfight processes than previous single-stage works. As an evaluation, we conduct experiments in the constructed dogfight environment. The experiment results demonstrate that the training environment constructed in this paper has high fidelity, effectively simulating the UCAV dogfight process and better aligning the physical model with real-world scenarios. The proposed learning framework performs excellently in the UCAV multi-dimensional decision-making task, significantly outperforming state-of-the-art deep RL and imitation learning algorithms, achieving higher learning efficiency, more effective air combat policies, and higher success rates. The contributions of this work can be summarized as follows.
@@ 2. Preliminaries
In this section, we review the preliminary knowledge about the Harfang3D Sandbox simulator, the Markov decision process, and the actor-critic framework.
@@ 2.1. Harfang3D Sandbox
We establish a simulation platform based on the Harfang3D sandbox (ozbek2022harfang3d, ), which is a customizable and flexible tool for combat aircraft dogfights. This platform facilitates the manipulation and access to flight dynamics models and environment states. The sandbox provides simulated aircraft of five different models, each equipped with unique missile configurations, and a limited missile supply. Therefore, the aircraft must lock onto a target before launching a missile to avoid missing the target and consequently losing a missile, which increases the task difficulty. Within this simulation environment, information about environment states can be obtained, including aircraft position, attitude, heading angle, and health as well as target lock status, thrust, lock angle, and missile loading. In particular, the sandbox also offers various unmanned aircraft control methods for components including rudders, elevators, ailerons, missile launch, engine thrust, landing gear, and flaps. As illustrated in Figure 1, rudders control the aircraft heading, elevators control the aircraft pitch, and ailerons control the aircraft roll. These three controls enable basic transformations of the aircraft’s attitude.
@@ 2.2. Markov Decision Process
The Markov Decision Process (MDP) (garcia2013markov, ) is the theoretical foundation of RL, which provides a formal description of the learning environment of the agent. In this paper, we model the UCAV dogfight environment as an MDP. The MDP is represented as a tuple $(S,A,P,R,\gamma)$, where $S$ is the state space of the environment, $A$ is the action space of the agent, $P$ is the state transition probability, with $P(s_{t+1}|s_{t},a_{t})$ indicating the probability of transitioning from state $s_{t}$ to state $s_{t+1}$ when taking action $a_{t}$ at time $t$, $R$ is the reward function, representing the immediate reward obtained by the intelligent agent after taking action $a$, and $\gamma$ is the reward discount factor, with $\gamma\in[0,1)$ and a $\gamma$ value closer to 1 indicating that the intelligent agent focuses more on long-term rewards. Reinforcement learning aims to learn policy function $\pi:S\rightarrow A$ that maximizes discounted reward $\mathbb{E}_{P,\pi}[\sum_{t=1}^{T}\gamma^{t}R(s_{t},a_{t})]$, where $T$ is the total timestep in an episode.
@@ 2.3. Actor-Critic Framework
The actor–critic framework (konda1999actor, ) is a policy optimization paradigm in RL that simultaneously optimizes both the value and policy functions, exhibiting good learning performance in sequential decision problems with continuous action spaces. The actor represents the policy $\pi$ of the agent, which specifies the actions to take at each state. It aims to learn an optimal policy that maximizes the expected cumulative reward. During training, the actor adjusts its policy based on feedback from the critic, gradually improving the policy performance. The critic estimates the value of each action to assess the quality of the policy and assists the actor in refining its policy. The actor–critic framework is based on policy gradient methods, combining back-propagation with value function methods to enhance the policy performance. Based on the actor-critic framework, the agent can learn a policy and corresponding value function during interaction with the environment, enabling high-level decision-making and intelligent behavior.
@@ 3. Related Work
With the rapid development of deep learning, the control for UCAVs has become more autonomous. Deep learning enables UCAVs to both execute fixed programs and exhibit learning capabilities. In recent years, deep learning techniques such as supervised learning, imitation learning, and RL have been widely used to address tactical decision-making problems in UCAV dogfights. In this section, we discuss the related works from four views: supervised learning methods, imitation learning methods, reinforcement learning methods, and methods of combining RL and imitation learning.
3.1. Supervised Learning
@@
Supervised learning is predominantly used in UCAV dogfights to predict the states after executing actions, providing a basis for action selection. Zhang and Huang (zhang2020maneuver, ) proposed a deep learning method integrating state evaluation and action selection for making flight action decisions. The state evaluation function scores the next state after executing an action and selects the action that generates the highest score. This approach enables autonomous flight control of UCAVs in one-on-one aerial combat. Li et al. (li2022decision, ) trained a long short-term memory network with time-series data, mapping historical state sequences onto flight control parameters. They used a stacked sparse autoencoder to reduce data redundancy and accelerate decision-making.
3.2. Imitation Learning
@@
Imitation learning (hussein2017imitation, ) is a machine learning paradigm where an agent tries to learn a policy by mimicking expert behavior. Imitation learning methods include 1) behavior cloning (BC), 2) inverse reinforcement learning, and 3) generative adversarial imitation learning.
@@
(1) BC (ly2020learning, ) involves training a model on expert demonstration data to resemble the expert behaviors. Sandström et al. (sandstrom2022fighter, ) adopted BC using control signals from human pilots to train a neural network, enabling the model to mimic the behaviors of human pilots for aircraft flight.
@@
(2) Inverse reinforcement learning (arora2021survey, ) estimates the reward function from expert demonstrations and uses it to guide the training of intelligent agents in reinforcement learning. By extracting the hidden reward function from expert data, the system can understand task objectives from expert demonstrations. Kong et al. (kong2020uav, ) employ a reward shaping method to address the problem of sparse rewards in reinforcement learning and use maximum entropy inverse reinforcement learning to obtain a shaped reward function, thereby accelerating convergence in training.
@@
(3) Generative adversarial imitation learning (ho2016generative, ) uses adversarial generative networks to learn hidden policies from expert demonstrations, imitating expert actions. Wang and Wei (wang2022research, ) applied this technique to address the problem of sparse rewards in reinforcement learning. They designed internal rewards to stimulate intelligent agents to explore autonomously, thus expediting training.
3.3. Reinforcement Learning